OpenStudy (anonymous):

Find the product of 5(cos pi/3 + i sin pi/3) x 3(cos pi + i sin pi)

OpenStudy (sirm3d):

\[\large 15(\cos \frac{ 4\pi }{ 3 }+i \sin \frac{ 4\pi }{ 3 })\]is this the result you are expecting or do you need a simplified form?

OpenStudy (anonymous):

no i got it